Each of the 50 students in class belongs to exactly one of the four groups A, B, C, or D. The membership numbers for the four groups are as follows: A: 5, B: 10, C: 15, D: 20. First, choose one of the 50 students at random and let X be the size of that student’s group. Next, choose one of the four groups at random and let Y be its size. (Recall: all random choices are with equal probability, unless otherwise specified.)

(a) Write down the probability mass functions for X and Y . (b) Compute E[X] and E[Y ] . (c) Compute V (X) and V (Y ). (d) Assume you have s students divided into n groups with membership numbers s1,s2,··· ,sn and again X is the size of the group of a randomly chosen student, while Y is the size of the randomly chosen group. Let E[Y ] = µ and V (Y ) = σ2. Express E[X] with s,n,µ, and σ.



1
Expert's answer
2022-04-18T00:20:57-0400

"a)\space P(X=5)=\frac{5}{5+10+15+20}=0.1\\
P(X=10)=\frac{10}{5+10+15+20}=0.2\\
P(X=15)=\frac{15}{5+10+15+20}=0.3\\
P(X=20)=\frac{20}{5+10+15+20}=0.4"

For all other values of X P(X) = 0

Probability of selecting one group out of 4 is 1 / 4 = 0.25

"P(Y=5)=0.25\\
P(Y=10)=0.25\\
P(Y=15)=0.25\\
P(Y=20)=0.25"

For all other values of Y P(Y) = 0


"b)\space\Epsilon[X]=0.1\cdot5+0.2\cdot10+\\
+0.3\cdot15+0.4\cdot20=15\\
\Epsilon[Y]=0.25\cdot(5+10+15+20)=12.5"


"c)\space V(X)=\frac{1}{4-1}(0.1\cdot(5-15)^2+0.2\cdot(10-15)^2+\\
0.3\cdot(15-15)^2+0.4\cdot(20-15)^2)]\approx\\
\approx8.333\\
V(Y)=\frac{1}{3}\cdot0.25\cdot(5+10+15+20)\approx4.167"


"d)\space \Epsilon[X]=\sum_{k=1}^{n}\frac{s_k}{s}s_k=\frac{1}{s}\sum_{k=1}^{n}s_k^2\\
\Epsilon[Y]=\sum_{k=1}^{n}\frac{1}{n}s_k=\frac{1}{n}\sum_{k=1}^{n}s_k=\frac{s}{n}\\
V[Y]=\frac{1}{n-1}\sum_{k=1}^{n}\frac{1}{n}(s_k-\frac{s}{n})^2=\\
=\frac{1}{n(n-1)}\sum_{k=1}^{n}(s_k^2-\frac{2s_ks}{n}+\frac{s^2}{n^2})=\\
=\frac{1}{n(n-1)}(s\Epsilon[X]-\frac{2s^2}{n}+\frac{s^2}{n})=\\
=\frac{1}{n(n-1)}(s\Epsilon[X]-\frac{s^2}{n})=\sigma^2\Rarr\\
\Rarr\Epsilon[X]=\frac{1}{s}(n(n-1)\sigma^2+\frac{s^2}{n})=\\
=\frac{n(n-1)}{s}\sigma^2+\mu"